OpenStudy (anonymous):

John Lawson owner of Lawson’s Lawns has been hired to mow the lawn at the Credit Union. John can complete the job by himself in 10 hours and his brother Tim can complete the job in 8 hours. How long would it take them to complete the job together?

OpenStudy (shane_b):

This is a related rate problem so think of it this way: rate1 + rate1 = combined rate: \[\frac{1}{10}+\frac{1}{8}=\frac{18}{80}\]To find the hours for 1 job you just do: \[\frac{1}{\frac{18}{80}}=4.44~hours\]